OpenStudy (anonymous):

a 5000W heater i used to heat water. How long will it take to heat 20kg water from 20C to 100C?

OpenStudy (mathmate):

Energy required =20*1000 g (100-20)*4.184 joules =6694400 joules Time required =6694400/5000 =1339 seconds =22min. 19 seconds